OpenStudy (anonymous):

4 - x = | 2 - 3x |, just confused with the absolute values! The | is meant to be the absolute value sign btw

OpenStudy (anonymous):

when someting is in absolute value it means it must be positive

OpenStudy (anonymous):

4-x=2-3x and 4-x=-(2-3x)

OpenStudy (anonymous):

solve for both situations

OpenStudy (anonymous):

x= -1

hartnn (hartnn):

when |a|=b a=b OR a=-b apply same thing here. u get 2 equations and 2 values of x

OpenStudy (anonymous):

okay! that's basically what i need, thank you!

hartnn (hartnn):

welcome ^_^